Assume a Company Xs stock price today and cost of equity is consistent with the dividend discount model for constant growth; HINT: Cost of Equity Dividend Yield + Capital Gain Yield. Use the information in the table below to answer problems 16-17. Any missing information will need to be solved for. TIME (Years) 0 NI 1,350.00 Dividends Shares Outstanding Book Value Price Per Share Return On Equity Cost of Equity Long Term Growth Dividends Per Share 135.00 $6,750.00 80.00 12.0% 16. Company Xs dividend payout ratio is closest to: A. 0.3 B. 0.4 C. 0.5 D. 0.6 E. 0.7 17. The present value of Company Xs first dividend payment one year from now (T1) is closest to: A. $3.33 B. $3.40 C. $3.73 D. $3.81 E. $3.83
Answer #1

Question 16

Return on equity = Net Income /Total equity (book value)

Return on Equity =ROE = 1350/6750 = 0.2= 20%

g = ROE * Retention ratio

g = 12% (given)

ROE = 20%

Retention Ratio = g /ROE = 0.12/0.20 = 0.6

Payout ratio = 1 - retention ratio = 1-0.6 = 0.4

Payout ratio = 0.4 (Option B)

Question 17:

Current Net income per share = 1350/135 =$10

Current dividend = $10*0.4 = $4

Dividend one year from now = Current dividend *(1+g) = 4* (1+0.12) = 4.48

Present value of the first dividend payment one year from now = 4.48/(1+r) = 4.48/(1+0.20) = 4.48/1.20 = 3.73

Present value of the first dividend payment one year from now= $3.73 (Option C)